OpenStudy (nickymail48):

0.1x+17.1=x-2.43 how to solve

OpenStudy (anonymous):

substract 0.1x both sides first

OpenStudy (nickymail48):

i did

OpenStudy (anonymous):

alright which gave you?

OpenStudy (nickymail48):

0.1x-2.43?

OpenStudy (anonymous):

you can't substract/add a number with a variable with only a number

OpenStudy (anonymous):

0.1x+17.1=x-2.43 -0.1x -0.1x

OpenStudy (nickymail48):

0.2x on rite side?

OpenStudy (anonymous):

x=1 so 1-0.1?

OpenStudy (nickymail48):

so its just 1

OpenStudy (anonymous):

No, 1-0.1 or 0.10

OpenStudy (nickymail48):

ok so its -9

OpenStudy (anonymous):

how come is negative if 1 is greater?

OpenStudy (anonymous):

and it's 0.9

OpenStudy (nickymail48):

0.9=2.43-17.01

OpenStudy (anonymous):

17.1=0.9x-2.43 +2.43= +2.43

OpenStudy (nickymail48):

19.53=0.9x

OpenStudy (anonymous):

right now divide 0.9 both sides and you get the answer :)

OpenStudy (nickymail48):

21.7

OpenStudy (nickymail48):

ok thanks!